# Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation

### v1

@@ B.1 Real-World Experiment Setup
B.2 Data Collection
We use the GELLO human teleoperate system [wu2024gello] to collect the data for real-world experiments. The GELLO controller device and the GELLO data collection system we use is illustrated in Figure 10. We perform hand-eye calibration in the real-world data collection. The calibration process enabled us to precisely determine the transformation relation from the camera coordinate system to the robot-base coordinate system. The observations including 3D point clouds and robot states and actions are recorded and consist the expert demonstration dataset. We use 3D point clouds as the robot’s visual input, which is directly acquired by the ZED-2i camera, a binocular camera can generate the point cloud of the scene. An example of the point cloud acquired by the camera is shown in Figure 11. A single point cloud output from the ZED-2i camera contains 252,672 points. The point cloud input to the policies including BridePolicy and baselines was further downsampled to 2048 points using the Farthest Point Sampling (FPS) algorithm and done by the fpsample library for further acceleration to support the real-time inference.
@@ B.2 Data Collection
Table 5: Evaluation of FlowPolicy using 10 step and 1 step when sampling
@@
Appendix C Model Implementation of BridgePolicy and Baselines
C.1 Implementation of The Denoising Network
Following the implementation of the denoising model in DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy], we adopt the U-net architecture as the backbone of the denoising model in BridgePolicy and the baselines. The numbers of parameters in the denoising model in BridgePolicy and the baselines are the same which is approximately 255M. The BridgePolicy possess additional modules, the semantic aligner and the modality fusion module, which only have about less than 1M parameters.
Appendix D Training and Inference Details of BridgePolicy and Baselines
D.1 Training
For the configurations such as the horizon length, the number of observation steps and the predicted action steps, we set them to the default setting in their publicly available code base and the corresponding configurations of BridgePolicy is aligned with DP3. Specifically, we list the hyperparameters used in BridgePolicy in Table 6 (Real-World).
@@ D.1 Training
basketball box-close coffee-pull coffee-push dial-turn hammer handle-pull pick-place-wall
0.1 1.00 0.68 0.95 0.94 0.92 1.00 0.59 0.98 0.88
0.3 1.00 0.72 0.95 0.93 0.91 0.97 0.53 0.94 0.87
0.5 1.00 0.75 0.88 0.93 0.87 0.97 0.56 0.93 0.86
0.6 1.00 0.75 0.96 0.90 0.93 0.97 0.54 0.91 0.87
0.8 1.00 0.81 0.96 0.83 0.85 0.98 0.53 0.93 0.86
1.0 1.00 0.66 0.92 0.92 0.85 0.99 0.57 0.93 0.85
1.2 1.00 0.86 0.91 0.91 0.78 0.96 0.56 0.93 0.86
1.5 1.00 0.75 0.92 0.93 0.78 0.96 0.56 0.88 0.85
D.2 Inference
Number of Sampling Steps. We set the number of sampling steps to be 10 for BridgePolicy and baselines except the number of sampling steps of FlowPolicy is set to 1 according to the original paper [zhang2025flowpolicy]. We also conduct additional experiments to explore how the performance would change if we increase the number of sampling steps and the results are shown in Table 5. These results show that the performance degrades, possibly due to the error accumulation issue in consistency flow matching [sabour2025align].
Appendix E Additional Experimental Results
We provide additional experimental results in this section. Despite the more ablation results and success rates of individual tasks, we also provide additional visualization of the real-world tasks shown in Figure 9 and 8.
E.1 More Ablation Results on $\alpha$ parameter
We provide more ablation results on the $\alpha$ parameter shown in Table 7 and 8. For most of the evaluated tasks, a smaller $\alpha$ less than $1$ is better for training the policy.
@@ E.1 More Ablation Results on $\alpha$ parameter
0.1 1.00 0.84 0.60 0.81 0.87 0.41 0.73 0.35 0.59
0.3 1.00 0.79 0.67 0.82 0.88 0.43 0.75 0.33 0.60
0.5 1.00 0.85 0.56 0.80 0.83 0.45 0.71 0.33 0.58
0.6 1.00 0.84 0.52 0.78 0.88 0.45 0.76 0.30 0.60
0.8 1.00 0.81 0.62 0.81 0.87 0.42 0.74 0.37 0.60
1.0 1.00 0.88 0.58 0.82 0.89 0.42 0.73 0.31 0.59
1.2 1.00 0.80 0.61 0.80 0.85 0.44 0.73 0.30 0.58
1.5 1.00 0.81 0.60 0.80 0.85 0.42 0.73 0.31 0.57
E.2 Parameter Sensitivity on $\gamma$ and $\lambda$
We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate the combinations of two hyperparameters (the steady variance level $\lambda^{2}$ (over 255) and the penalty coefficient $\gamma$) in the UniDB [zhu2025unidb] framework and the results are shown in Table 9. Results in Table 9 demonstrates that $\lambda=50$ and $\gamma=10^{7}$ are the relatively better hyperparameter combination of the UniDB framework in policy learning.
@@ E.2 Parameter Sensitivity on $\gamma$ and $\lambda$
E.3 Success Rates of Individual Tasks
@@
DP [chi2023diffusion] 0.45±0.05 0.35±0.04 0.15±0.04 0.69±0.08 0.23±0.09 0.63±0.08 0.23±0.08
DP3 [ze20243d] 1.00±0.00 0.65±0.05 0.40±0.05 0.86±0.08 0.36±0.04 0.73±0.07 0.32±0.07
Simple DP3 [ze20243d] 1.00±0.00 0.59±0.01 0.45±0.03 0.79±0.07 0.26±0.04 0.63±0.07 0.22±0.07
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.58±0.06 0.53±0.12 0.80±0.06 0.38±0.09 0.70±0.06 0.28±0.05
@@
DP [chi2023diffusion] 0.93±0.01 0.98±0.01 0.96±0.03 0.93±0.03 0.99±0.01 0.63±0.10 1.00±0.00
DP3 [ze20243d] 1.00±0.00 1.00±0.00 0.99±0.02 0.99±0.01 1.00±0.00 0.66±0.01 1.00±0.00
Simple DP3 [ze20243d] 1.00±0.00 1.00±0.00 0.98±0.02 0.93±0.02 1.00±0.00 0.59±0.01 0.97±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.97±0.03 0.98±0.02 1.00±0.00 1.00±0.00 0.88±0.10 0.90±0.08
@@
DP [chi2023diffusion] 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 [ze20243d] 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 [ze20243d] 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
@@
DP [chi2023diffusion] 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 [ze20243d] 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 [ze20243d] 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
@@
DP [chi2023diffusion] 0.99±0.00 0.75±0.04 1.00±0.00 0.85±0.06 0.15±0.04 0.30±0.05 0.15±0.06
DP3 [ze20243d] 1.00±0.00 1.00±0.01 1.00±0.00 0.98±0.02 0.38±0.30 0.42±0.03 0.72±0.04
Simple DP3 [ze20243d] 0.99±0.00 1.00±0.01 1.00±0.00 0.95±0.04 0.28±0.16 0.38±0.05 0.62±0.09
FlowPolicy [zhang2025flowpolicy] 0.65±0.05 1.00±0.00 1.00±0.00 0.66±0.06 0.66±0.14 0.81±0.04 0.98±0.02
@@
DP [chi2023diffusion] 0.34±0.07 0.20±0.03 0.14±0.04 0.34±0.07 0.67±0.04 0.18±0.08 0.10±0.04
DP3 [ze20243d] 0.67±0.07 0.51±0.08 0.18±0.03 0.85±0.03 0.94±0.03 0.96±0.03 0.17±0.05
Simple DP3 [ze20243d] 0.48±0.07 0.38±0.08 0.16±0.03 0.92±0.12 0.86±0.06 0.88±0.03 0.09±0.05
FlowPolicy [zhang2025flowpolicy] 0.70±0.09 0.63±0.08 0.33±0.05 0.96±0.02 0.61±0.06 0.70±0.04 0.31±0.02
@@
DP [chi2023diffusion] 0.00±0.00 0.02±0.01 0.15±0.01 0.28±0.03 0.09±0.02
DP3 [ze20243d] 0.14±0.09 0.12±0.04 0.99±0.01 0.62±0.03 0.14±0.04
Simple DP3 [ze20243d] 0.08±0.06 0.12±0.06 0.79±0.01 0.32±0.03 0.12±0.05
FlowPolicy [zhang2025flowpolicy] 0.31±0.04 0.63±0.05 1.00±0.00 0.73±0.05 0.26±0.02
@@
DP [chi2023diffusion] 0.63±0.03 0.11±0.02 0.11±0.03 0.05±0.01 0.43±0.07
DP3 [ze20243d] 0.97±0.04 0.27±0.08 0.19±0.10 0.35±0.08 0.75±0.04
Simple DP3 [ze20243d] 0.97±0.05 0.15±0.08 0.05±0.01 0.28±0.05 0.50±0.03
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.56±0.03 0.40±0.03 0.95±0.03 0.88±0.03
@@
E.4 More Sample Efficiency

### v2

@@ C.2 Real-World Experiment Setup
C.3 Data Collection
We use the GELLO human teleoperate system (Wu et al., 2024a) to collect the data for real-world experiments. The GELLO controller device and the GELLO data collection system we use is illustrated in Figure 7. We perform hand-eye calibration in the real-world data collection. The calibration process enabled us to precisely determine the transformation relation from the camera coordinate system to the robot-base coordinate system. The observations including 3D point clouds and robot states and actions are recorded and consist the expert demonstration dataset. We use 3D point clouds as the robot’s visual input, which is directly acquired by the ZED-2i camera, a binocular camera can generate the point cloud of the scene. An example of the point cloud acquired by the camera is shown in Figure 8. A single point cloud output from the ZED-2i camera contains 252,672 points. The point cloud input to the policies including BridePolicy and baselines was further downsampled to 2048 points using the Farthest Point Sampling (FPS) algorithm and done by the fpsample library (Han et al., 2023) for further acceleration to support the real-time inference.
@@ C.3 Data Collection
Table 6: Evaluation of FlowPolicy using 10 step and 1 step for sampling.
@@
Appendix D Model Implementation of BridgePolicy and Baselines
D.1 Implementation of The Denoising Network
Following the implementation of the denoising model in DP3 (Ze et al., 2024) and FlowPolicy (Zhang et al., 2025), we adopt the U-Net architecture as the backbone of the denoising model in BridgePolicy and the baselines. The numbers of parameters in the denoising model in BridgePolicy and the baselines are the same which is approximately 255M. The BridgePolicy possess additional modules, the semantic aligner and the modality fusion module, which only have about less than 1M parameters.
Appendix E Training and Inference Details of BridgePolicy and Baselines
E.1 Training
For the configurations such as the horizon length, the number of observation steps and the predicted action steps, we set them to the default setting in their publicly available code base and the corresponding configurations of BridgePolicy is aligned with DP3. For training of the aligner in BridgePolicy, we frozen the parameters after 50 epochs. Specifically, we list the hyperparameters used in BridgePolicy in Table 7 (Real-World).
@@ E.1 Training
coffee-push hammer pick-place-wall
0.0 0.91 0.96 0.88 0.91
0.5 0.99 0.97 0.93 0.96
1.0 0.91 0.99 0.96 0.95
2.0 0.91 0.97 0.87 0.91
5.0 0.92 0.93 0.83 0.89
E.2 Inference
Number of Sampling Steps. We set the number of sampling steps to be 10 for BridgePolicy and baselines except the number of sampling steps of FlowPolicy is set to 1 according to the original paper (Zhang et al., 2025). We also conduct additional experiments to explore how the performance would change if we increase the number of sampling steps and the results are shown in Table 6. These results show that the performance degrades, possibly due to the error accumulation issue in consistency flow matching (Sabour et al., 2025).
Appendix F Additional Experimental Results
Here we provide additional experimental results.
F.1 Additional Visualization of Real-World Tasks
Figure 10, 11, and 12 demonstrate the additional visualization of the real-world tasks between our BridgePolicy and baselines.
F.2 More Ablation Results on $\alpha$ parameter
We provide more ablation results on the $\alpha$ parameter shown in Table 8 and 9. For most of the evaluated tasks, a smaller $\alpha$ less than $1$ is better for training the policy.
@@ F.2 More Ablation Results on $\alpha$ parameter
0.0 1.00 0.79 0.56 0.78 0.85 0.42 0.74 0.34 0.58
0.5 1.00 0.83 0.51 0.78 0.86 0.41 0.75 0.35 0.59
1.0 1.00 0.84 0.54 0.79 0.91 0.44 0.76 0.30 0.60
2.0 1.00 0.84 0.62 0.82 0.86 0.42 0.76 0.32 0.59
5.0 0.97 0.80 0.57 0.78 0.89 0.43 0.73 0.30 0.58
F.3 Parameter Sensitivity on $\gamma$ and $\lambda$
We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate the combinations of two hyperparameters (the steady variance level $\lambda^{2}$ (over 255) and the penalty coefficient $\gamma$) in the UniDB (Zhu et al., 2025) framework and the results are shown in Table 10. Results in Table 10 demonstrates that $\lambda=50$ and $\gamma=10^{7}$ are the relatively better hyperparameter combination of the UniDB framework in policy learning.
@@ F.3 Parameter Sensitivity on $\gamma$ and $\lambda$
F.4 Success Rates of Individual Tasks
@@
DP 0.45±0.05 0.35±0.04 0.15±0.04 0.69±0.08 0.23±0.09 0.63±0.08 0.23±0.08
DP3 1.00±0.00 0.65±0.05 0.40±0.05 0.86±0.08 0.36±0.04 0.73±0.07 0.32±0.07
Simple DP3 1.00±0.00 0.59±0.01 0.45±0.03 0.79±0.07 0.26±0.04 0.63±0.07 0.22±0.07
FlowPolicy 1.00±0.00 0.58±0.06 0.53±0.12 0.80±0.06 0.38±0.09 0.70±0.06 0.28±0.05
VITA 0.96±0.02 0.81±0.04 0.55±0.05 0.82±0.05 0.39±0.04 0.72±0.05 0.28±0.06
@@
DP 0.93±0.01 0.98±0.01 0.96±0.03 0.93±0.03 0.99±0.01 0.63±0.10 1.00±0.00
DP3 1.00±0.00 1.00±0.00 0.99±0.02 0.99±0.01 1.00±0.00 0.66±0.01 1.00±0.00
Simple DP3 1.00±0.00 1.00±0.00 0.98±0.02 0.93±0.02 1.00±0.00 0.59±0.01 0.97±0.00
FlowPolicy 1.00±0.00 0.97±0.03 0.98±0.02 1.00±0.00 1.00±0.00 0.88±0.10 0.90±0.08
VITA 0.99±0.01 1.00±0.00 0.99±0.01 1.00±0.00 1.00±0.00 0.90±0.20 1.00±0.00
@@
DP 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
VITA 0.90±0.01 0.95±0.01 0.68±0.07 1.00±0.00 0.19±0.00 0.76±0.05 0.73±0.12
@@
DP 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
VITA 0.34±0.10 0.54±0.10 0.99±0.01 0.81±0.12 1.00±0.00 0.71±0.07 0.99±0.01
@@
DP 0.99±0.00 0.75±0.04 1.00±0.00 0.85±0.06 0.15±0.04 0.30±0.05 0.15±0.06
DP3 1.00±0.00 1.00±0.01 1.00±0.00 0.98±0.02 0.38±0.30 0.42±0.03 0.72±0.04
Simple DP3 0.99±0.00 1.00±0.01 1.00±0.00 0.95±0.04 0.28±0.16 0.38±0.05 0.62±0.09
FlowPolicy 0.65±0.05 1.00±0.00 1.00±0.00 0.66±0.06 0.66±0.14 0.81±0.04 0.98±0.02
VITA 1.00±0.00 0.96±0.03 1.00±0.00 0.90±0.01 0.23±0.01 0.52±0.06 0.90±0.01
@@
DP 0.34±0.07 0.20±0.03 0.14±0.04 0.34±0.07 0.67±0.04 0.18±0.08 0.10±0.04
DP3 0.67±0.07 0.51±0.08 0.18±0.03 0.85±0.03 0.94±0.03 0.96±0.03 0.17±0.05
Simple DP3 0.48±0.07 0.38±0.08 0.16±0.03 0.92±0.12 0.86±0.06 0.88±0.03 0.09±0.05
FlowPolicy 0.70±0.09 0.63±0.08 0.33±0.05 0.96±0.02 0.61±0.06 0.70±0.04 0.31±0.02
VITA 0.42±0.11 0.51±0.16 0.19±0.04 1.00±0.00 0.63±0.06 0.87±0.04 0.12±0.01
@@
DP 0.00±0.00 0.02±0.01 0.15±0.01 0.28±0.03 0.09±0.02
DP3 0.14±0.09 0.12±0.04 0.99±0.01 0.62±0.03 0.14±0.04
Simple DP3 0.08±0.06 0.12±0.06 0.79±0.01 0.32±0.03 0.12±0.05
FlowPolicy 0.31±0.04 0.63±0.05 1.00±0.00 0.73±0.05 0.26±0.02
VITA 0.23±0.06 0.43±0.10 0.99±0.01 0.65±0.04 0.09±0.02
@@
DP 0.63±0.03 0.11±0.02 0.11±0.03 0.05±0.01 0.43±0.07
DP3 0.97±0.04 0.27±0.08 0.19±0.10 0.35±0.08 0.75±0.04
Simple DP3 0.97±0.05 0.15±0.08 0.05±0.01 0.28±0.05 0.50±0.03
FlowPolicy 1.00±0.00 0.56±0.03 0.40±0.03 0.95±0.03 0.88±0.03
VITA 0.83±0.07 0.61±0.03 0.16±0.05 0.80±0.09 0.69±0.03
@@
F.5 More Sample Efficiency